# Fast Compression of Deterministic Finite Automata

### v2

@@ 1 Introduction
Processing a character D2FA during matching may require following multiple default transitions, thus incurring (as the name suggests) a delay. Minimizing the delay is essential in high-throughput applications, and two important variants of D2FAs that address have been proposed. Kumar et al. [27] gave a modified D2FA construction that, given an integer parameter $L$, limits the maximum path of default transitions to $L$. We refer to this as the longest delay variant of the problem. Alternatively, Becchi and Crowley [7, 9] gave a modified D2FA construction that limits the maximum total number of default transitions traversed on any input string $S$ by $|S|$. We refer to this as the matching delay variant.
@@ 2 Preliminaries
A graph $G=(V,E)$ is a set of nodes $V$ (also called vertices) and a set of edges $E:V\times V$ between nodes. We call the two nodes of an edge its endpoints. If the edges have direction, i.e. $(u,v)\neq(v,u)$, then the graph is directed, otherwise it is undirected. Edges can have an associated weight, in which case the graph is weighted. Edge can also have an associated label, in which case the graph is labeled. We denote an edge from $u$ to $v$ with label $c$ as ${(u,v)}_{c}$ and say $(u,v)$ is $c$-labeled. A path of length $k$ between two nodes $u_{0}$ and $u_{k}$ is a sequence of nodes $u_{0},\ldots,u_{k}$ such that $(u_{i},u_{i+1})\in E$ for $0\leq i<k$. It can also be viewed as the sequence of edges $(u_{0},u_{1}),\ldots,(u_{k-1},u_{k})$. If $u_{0}=u_{k}$ then $p$ is a cycle. A set of nodes where every pair of nodes have a path between them is called a connected component. If a graph has only one connected component it is connected, otherwise it is disconnected.
@@ 3 Delayed Deterministic Finite Automata
A delayed deterministic finite automaton (D2FA) [27] is a deterministic finite automaton that is augmented with unlabeled default transitions. Formally a D2FA is a 6-tuple $D^{2}=(Q,\Sigma,\delta,q_{0},A,F)$. As for DFAs, $Q$ is the set of states, $\Sigma$ is the alphabet, $\delta$ is the transition functions, $q_{0}\in Q$ is the initial state, and $A\subseteq Q$ is the set of accepting states. The final component $F:Q\rightarrow Q$ is the default transition function. Viewed as a graph, default transitions are $\epsilon$-labeled directed edges, where $\epsilon$ is the empty string, and each state has at most one outgoing default transition. See Figure 1 (C) for an example. To transition from a state $u$ according to a character $c$, we follow a $c$-labeled transition if it exists or otherwise follow the default transition:
$\delta(u,c)=\begin{cases}v&\text{ if }(u,v)_{c}\in D^{2}\\ \delta(F(u),c)&\text{ otherwise }\end{cases}$
@@
Given a character $c$ and a path $p=u_{1},\ldots,u_{k}$ we say $p$ matches $c$ if $\delta(u_{1},c)=u_{k}$ and all but the last transition is default, i.e., $F(u_{i})=u_{i+1}$ for $1\leq i<k$. Note that the concatenation of the labels of $p$ equals $c$. Given a string $S=c_{1}\ldots c_{m}$ and a path $p$ we say $p$ matches $S$ if $p$ is the concatenation of the paths matching the individual characters $c_{1},\ldots,c_{m}$. Note that the concatenation of the labels of $p$ is $S$. We define acceptance as before. Two D2FAs are equivalent if they have the same language, and two transitions are equivalent if they have the same destination and label.
@@
[…]
@@ 4 Compressing DFAs with Default Transitions
[…]
@@
Step 1 uses $O(n^{2}|\Sigma|$ to compare all states and construct the SRG. Step 2 constructs the maximum spanning tree with Kruskal’s algorithm [25] in $O(n^{2}\log n)$ time, and finally step 3 takes $O(n|\Sigma|)$ time. In total, the running time is $O(n^{2}\log n+n^{2}|\Sigma|))$ [^1] The running time is not explicitly stated in the paper, but follows from the description..
@@ 5 Fast Compression
Let $D$ be the input DFA and let $r$ and $k$ be two constant, positive integer parameters. We initialize an undirected graph $G=(Q,E)$ where $Q$ is the set of states in $D$ and $E=\{(q_{0},u)\ |\ u\neq v_{0}\}$, i.e., we have a edge between the initial state and all other states. We then add edges to the graph in $r$ rounds, where each round proceeds as follows:
@@
The above scheme is inspired the locality-sensitive scheme by Har-Peled et al. [22] for bitstrings w.r.t. Hamming distance. Their solution sample positions from the input bitstrings. In our solution, the sampled positions correspond to the sampled characters from $\Sigma$.
@@ 6.2 Fast Compression with Bounded Longest Delay
[…]
@@
[…]
@@ 7.1 Bounded Matching Delay by the A-DFA Algorithm
Initialize a D2FA $D^{2}$ no default transitions. We proceed as follows.
@@
[…]
@@ 7.2 Fast Compression with Bounded Matching Delay
[…]
Hashing a state and computing the similarity of the potential new default transition uses $O(k|\Sigma|)$ time. Hence, the full algorithm uses $O(rkn+rn|\Sigma|)=O(n\Sigma)$ time.
@@ 8 Experimental Evaluation
We implemented our methods described in the previous section and measured their performance on regular expressions extracted from widely used intrusion detection systems.
@@ 8.4 Results
Figure 2: Results for the Snort dataset on the algorithms for general compression (top), bounded longest delay (middle), and bounded longest matching delay (bottom). On the left, we show compression time in seconds vs. the number of states in the input DFA. On the right, we show the number of states in D2FA as a percent of the number of states in the input DFA.
Figure 3: Results for the Suricata dataset on the algorithms for general compression (top), bounded longest delay (middle), and bounded longest matching delay (bottom). On the left, we show compression time in seconds vs. the number of states in the input DFA. On the right, we show the number of states in D2FA as a percent of the number of states in the input DFA.
Figure 4: Results for the Zeek dataset on the algorithms for general compression (top), bounded longest delay (middle), and bounded longest matching delay (bottom). On the left, we show compression time in seconds vs. the number of states in the input DFA. On the right, we show the number of states in D2FA as a percent of the number of states in the input DFA.
@@ 9 Conclusion
An interesting open problem is to explore if our new framework can be combined with other DFA compression techniques, such as one mentioned in Section 1.2. Substantial work has been done on compression DFAs and we believe it is likely that combinations with our technique will lead to even faster solutions.
@@ Appendix A Input DFA Characteristics
Source
States
Rules
Average length of rules
% using wildcards (*, +, ?)
% using length restrictions ({,k,+})
Snort
991
12
34.3
50.0
33.3
Snort
1691
14
35.2
57.1
42.9
Snort
5667
15
37.3
60.0
46.7
Snort
8467
22
35.3
59.1
40.9
Snort
15902
50
29.3
46.0
30.0
Snort
39244
61
28.5
41.0
24.6
Snort
83055
101
28.0
53.5
15.8
Snort
144413
102
28.3
53.9
16.7
Snort
246547
128
29.9
49.2
23.4
Snort
523737
134
30.0
48.5
25.4
Snort
1069239
175
37.9
55.4
39.4
Suricata
1066
25
101.6
92.0
0.0
Suricata
2691
28
106.2
89.3
0.0
Suricata
6211
30
112.8
90.0
0.0
Suricata
13577
32
117.2
90.6
0.0
Suricata
32014
38
111.1
86.8
2.6
Suricata
75596
49
105.7
85.7
2.0
Suricata
128696
93
103.6
87.1
1.1
Suricata
273904
101
103.0
87.1
1.0
Suricata
462248
180
101.0
89.4
0.6
Suricata
995152
194
99.1
87.1
0.5
Zeek
2338
5
33.2
60.0
40.0
Zeek
4265
13
46.0
61.5
15.4
Zeek
10137
14
46.2
64.3
14.3
Zeek
12747
17
40.7
52.9
11.8
Zeek
30131
18
42.3
55.6
11.1
Zeek
67933
19
43.7
57.9
10.5
Zeek
127839
20
43.6
60.0
10.0
Zeek
340279
21
44.6
61.9
9.5
Zeek
384949
27
39.6
55.6
7.4
Zeek
1036954
29
39.6
58.6
10.3

### v4

# Philip Bille

@@ 1 Introduction
Processing a character D2FA during matching may require following multiple default transitions, thus incurring (as the name suggests) a delay. Minimizing the delay is essential in high-throughput applications, and two important variants of D2FAs that address this problem have been proposed. Kumar et al. [29] gave a modified D2FA construction that, given an integer parameter $L$, limits the maximum path of default transitions to $L$. We refer to this as the longest delay variant of the problem. Alternatively, Becchi and Crowley [7, 9] gave a modified D2FA construction that limits the maximum total number of default transitions traversed on any input string $S$ by $|S|$. We refer to this as the matching delay variant.
@@ 2 Preliminaries
A graph $G=(V,E)$ is a set of nodes $V$ (also called vertices) and a set of edges $E:V\times V$ between nodes. We call the two nodes of an edge its endpoints. If the edges have direction, i.e., $(u,v)\neq(v,u)$, then the graph is directed, otherwise it is undirected. Edges can have an associated weight, in which case the graph is weighted. An edge can also have an associated label, in which case the graph is labeled. We denote an edge from $u$ to $v$ with label $c$ as ${(u,v)}_{c}$ and say $(u,v)$ is $c$-labeled. A path of length $k$ between two nodes $u_{0}$ and $u_{k}$ is a sequence of nodes $u_{0},\ldots,u_{k}$ such that $(u_{i},u_{i+1})\in E$ for $0\leq i<k$. It can also be viewed as the sequence of edges $(u_{0},u_{1}),\ldots,(u_{k-1},u_{k})$. If $u_{0}=u_{k}$ then $p$ is a cycle. A set of nodes where every pair of nodes has a path between them is called a connected component. If a graph has only one connected component, it is connected, otherwise it is disconnected.
@@ 3 Delayed Deterministic Finite Automata
A delayed deterministic finite automaton (D2FA) [29] is a deterministic finite automaton that is augmented with unlabeled default transitions. Formally a D2FA is a 6-tuple $D^{2}=(Q,\Sigma,\delta,q_{0},A,F)$. As for DFAs, $Q$ is the set of states, $\Sigma$ is the alphabet, $\delta$ is the transition functions, $q_{0}\in Q$ is the initial state, and $A\subseteq Q$ is the set of accepting states. The final component $F:Q\rightarrow Q$ is the default transition function. Viewed as a graph, default transitions are $\epsilon$-labeled directed edges, where $\epsilon$ is the empty string, and each state has at most one outgoing default transition. See Figure 1 (C) for an example.
To transition from a state $u$ according to a character $c$, we follow a $c$-labeled transition if it exists or otherwise follow the default transition:
$\delta(u,c)=\begin{cases}v&\text{ if }(u,v)_{c}\text{ is a edge.}\\ \delta(F(u),c)&\text{ otherwise }\end{cases}$
@@
Given a character $c$ and a path $p=u_{1},\ldots,u_{k}$ we say $p$ matches $c$ if $\delta(u_{1},c)=u_{k}$ and all but the last transition is default, i.e., $F(u_{i})=u_{i+1}$ for $1\leq i<k$. Note that the concatenation of the labels of $p$ equals $c$. Given a string $S=c_{1}\ldots c_{m}$ and a path $p$ we say $p$ matches $S$ if $p$ is the concatenation of paths matching the individual characters $c_{1},\ldots,c_{m}$. Note that the concatenation of the labels of $p$ is $S$. We define acceptance as before. Two D2FAs are equivalent if they have the same language, and two transitions are equivalent if they have the same destination and label.
@@
Given a DFA, we can compress it by replacing sets of equivalent transitions with single default transitions to obtain an equivalent D2FA with fewer total transitions. We define the similarity of two states $u$ and $v$, denoted $\textit{sim}(u,v)$, to be their number of equivalent transitions, that is, $\textit{sim}(u,v)=|\{c\in\Sigma\mid\delta(u,c)=\delta(v,c)\}|$. See Figure 1 (B) for an illustration. Inserting a default transition $(u,v)$ and removing the equivalent transitions from $u$ does not affect the language but saves $\textit{sim}(u,v)-1$ transitions. Each transition we can remove without affecting the language we say is redundant.
@@ 4 Compressing DFAs with Default Transitions
Construct the complete, undirected graph over the states of $D$, and to each edge $(u,v)$ assign weight $\textit{sim}(u,v)$. This is the space reduction graph (SRG). See Figure 1 (B) for an example.
@@
Step 1 uses $O(n^{2}|\Sigma|)$ to compare all states and construct the SRG. Step 2 constructs the maximum spanning tree with Kruskal’s algorithm [27] in $O(n^{2}\log n)$ time, and finally step 3 takes $O(n|\Sigma|)$ time. In total, the running time is $O(n^{2}\log n+n^{2}|\Sigma|))$ [^2] The running time is not explicitly stated in the paper, but follows from the description..
@@ 5 Fast Compression
Let $D$ be the input DFA and let $r$ and $k$ be two constant, positive integer parameters. We initialize an undirected graph $G=(Q,E)$ where $Q$ is the set of states in $D$ and $E=\{(q_{0},u)\ |\ u\neq q_{0}\}$, i.e., we have a edge between the initial state and all other states. We then add edges to the graph in $r$ rounds, where each round proceeds as follows:
@@
The above scheme is inspired by the locality-sensitive scheme by Har-Peled et al. [23] for bitstrings w.r.t. Hamming distance. Their solution samples positions from the input bitstrings. In our solution, the sampled positions correspond to the sampled characters from $\Sigma$.
@@ 6.2 Fast Compression with Bounded Longest Delay
Construct a maximum spanning tree $T_{0}$ over $G$ using Kruskal’s algorithm. Pick a central node $v_{0}$ in $T_{0}$ and then discard $T_{0}$. Construct a new maximum spanning tree $T$ using Prim’s algorithm [40] with $v_{0}$ as the initial node. When we queue a new edge $(u,v)$, where $u$ is the node already in $T$, assign weight $w^{\prime}_{u,v}=\textit{sim}(u,v)-2^{d_{v}}$ where $d_{v}$ is the distance from $v_{0}$ to $v$ in $T$.
@@
Each edge $(u,v)$ we cut results in $\textit{sim}(u,v)-1$ more labeled transitions in $D^{2}$, as that default transition is then not constructed. Intuitively, the lower the diameter of $T$, the fewer edges we cut to get each tree below the bound. Therefore, we use an edge weight that trades similarity for lower diameter, as we observed that the fewer cuts outweighed the lost similarity.
@@ 7.1 Bounded Matching Delay by the A-DFA Algorithm
Initialize a D2FA $D^{2}$ with no default transitions. We proceed as follows.
@@
For each state $u\in Q$ add default transition $(u,v)$ to $D^{2}$, where $v$ is the state such that $\textit{sim}(u,v)$ is maximum and $d(v)<d(u)$.
@@ 7.2 Fast Compression with Bounded Matching Delay
First, pick $k$ unique random characters $c_{1},\ldots,c_{k}\in\Sigma$. Then, for each state $v\in Q$ we construct the sequence $V=\delta(v,c_{1}),\ldots,\delta(v,c_{k})$, and hash $V$ into a single hash value $h(v)$. We insert $v$ into a table with key $h(v)$. For each unique hash value $h_{i}$, we consider the set of states $C_{i}$ that hash to that value. For each state $u\in C_{i}$ we pick another state $v\in C_{i}$ uniformly at random. If $v$ has lower depth and the default transition $(u,v)$ compresses better than the current default transition of $u$, i.e., $d(v)<d(u)$ and $\textit{sim}(u,v)>\textit{sim}(u,F(u))$, we update the default transition of $u$ to point to $v$ in $D^{2}$, otherwise, we continue. After $r$ rounds, the algorithm terminates and returns the resulting D2FA $D^{2}$.
Hashing a state and computing the similarity of the potential new default transition uses $O(k|\Sigma|)$ time. Hence, the full algorithm uses $O(rkn+rn|\Sigma|)=O(n|\Sigma|)$ time.
@@ 8 Experimental Evaluation
We implemented our methods described in the previous section and measured their performance on regular expressions extracted from widely used intrusion detection systems. The implementation is available at https://github.com/MaxRishoj/fcomp-dfa.
@@ 8.4 Results
Figure 2: Results for the Snort dataset on the algorithms for general compression (top), bounded longest delay (middle), and bounded longest matching delay (bottom). On the left, we show compression time in seconds vs. the number of states in the input DFA. On the right, we show the number of transitions in the D2FA as a percent of the number of transitions in the input DFA.
Figure 3: Results for the Suricata dataset on the algorithms for general compression (top), bounded longest delay (middle), and bounded longest matching delay (bottom). On the left, we show compression time in seconds vs. the number of states in the input DFA. On the right, we show the number of transitions in the D2FA as a percent of the number of transitions in the input DFA.
Figure 4: Results for the Zeek dataset on the algorithms for general compression (top), bounded longest delay (middle), and bounded longest matching delay (bottom). On the left, we show compression time in seconds vs. the number of states in the input DFA. On the right, we show the number of transitions in the D2FA as a percent of the number of transitions in the input DFA.
@@ 9 Conclusion
An interesting open problem is to explore if our new framework can be combined with other DFA compression techniques, such as the ones mentioned in Section 1.2. Substantial work has been done on compression DFAs, and we believe it is likely that combinations with our technique will lead to even faster solutions.
10 Acknowledgments
This paper is inspired by earlier work in an MSc. thesis [24] supervised by two of the authors. We thank the anonymous reviewers of earlier drafts of this article for many valuable comments that improved the quality of the work.
@@ Appendix A Input DFA Characteristics
Source States Rules Average length of rules % using wildcards (*, +, ?) % using length restrictions ({,k,+})
Snort 991 12 34.3 50.0 33.3
Snort 1691 14 35.2 57.1 42.9
Snort 5667 15 37.3 60.0 46.7
Snort 8467 22 35.3 59.1 40.9
Snort 15902 50 29.3 46.0 30.0
Snort 39244 61 28.5 41.0 24.6
Snort 83055 101 28.0 53.5 15.8
Snort 144413 102 28.3 53.9 16.7
Snort 246547 128 29.9 49.2 23.4
Snort 523737 134 30.0 48.5 25.4
Snort 1069239 175 37.9 55.4 39.4
Suricata 1066 25 101.6 92.0 0.0
Suricata 2691 28 106.2 89.3 0.0
Suricata 6211 30 112.8 90.0 0.0
Suricata 13577 32 117.2 90.6 0.0
Suricata 32014 38 111.1 86.8 2.6
Suricata 75596 49 105.7 85.7 2.0
Suricata 128696 93 103.6 87.1 1.1
Suricata 273904 101 103.0 87.1 1.0
Suricata 462248 180 101.0 89.4 0.6
Suricata 995152 194 99.1 87.1 0.5
Zeek 2338 5 33.2 60.0 40.0
Zeek 4265 13 46.0 61.5 15.4
Zeek 10137 14 46.2 64.3 14.3
Zeek 12747 17 40.7 52.9 11.8
Zeek 30131 18 42.3 55.6 11.1
Zeek 67933 19 43.7 57.9 10.5
Zeek 127839 20 43.6 60.0 10.0
Zeek 340279 21 44.6 61.9 9.5
Zeek 384949 27 39.6 55.6 7.4
Zeek 1036954 29 39.6 58.6 10.3